The following output summarizes the results of fitting a least-squares regression to simulated data. Because we constructed these data using a computer, we know the SRM holds and we know the parameters of the model. We chose Beta0 = 7, Beta1 = 0.5, and se = 1.5. The fit is to a sample of = 50 cases


· If Beta1 = ½ in the population, then why isn’t b1 = ½?

· Do the 95% confidence intervals for Beta0 and Beta1 contain Beta0 and Beta1 in this example?

· What’s going to change in this summary if we increase the sample size from 50 cases to 5,000 cases?











1
Expert's answer
2021-01-27T15:55:04-0500


a) The values of "\beta_1" and b1 are not equal because b1 is only an estimate of the population slope that will change due to sampling variation.

b) Degree of freedom = n-2 = 50-2 = 48

Critical value of t at 95% confidence interval and df = 48 is 2.01

95% confidence interval of "\beta_0" is

"(6.993459 -2.01 \times 0.181933, 6.993459 + 2.01 \times 0.181933) \\

(6.993459 -0.365685, 6.993459 + 0.365685) \\

(6.627774, 7.359144)"

The 95% confidence interval for "\beta_0" contains the value of "\beta_0"

95% confidence interval of "\beta_1" is

"(0.5134397 -2.01 \times 0.029887, 0.5134397 + 2.01 \times 0.029887) \\

( 0.5134397 -0.060072, 0.5134397 +0.060072) \\

(0.4533677, 0.5735117)"

The 95% confidence interval for "\beta_1" contains the value of "\beta_1"

c) If we increase the sample size, the model will overfit the data.

SSError will decrease and SSRregression will increase

So, the value of r^2 will increase.

The value of se will decrease.

The values of standard errors will decrease.